OpenStudy (anonymous):

2a+2 ----- a²-1 Help?

OpenStudy (anonymous):

\[\frac{ 2a+2 }{ a²-1 }\]

OpenStudy (anonymous):

start with \[\frac{2(a+1)}{(a+1)(a-1)}\] and then cancel

OpenStudy (anonymous):

leaves you with \[\frac{2}{a-1}\]

OpenStudy (anonymous):

Can u simlify more?

OpenStudy (anonymous):

not after you write \(\frac{2}{a-1}\) that is the best you can do

OpenStudy (e.mccormick):

Are they also asking for the restrictions on that? If so, that is when the division would be undefined. Does that make sense?

OpenStudy (anonymous):

Oh, ok. Thanks!!!